1. Jess is 12 years older than steph. Ten years ago, Jess was two times older than steph. How old are they now ?


2.Bart is 16 years older than
Lisa. In three years, Bart will be three times as old as lisa. How old are they now ?

Answer by ewatrrr(24785)


Hi
1. Jess is 12 years older than steph
Let x and (x+12)represent steph's and jess' ages NOW respectively
Question states*** 10 years ago...
(x+12) - 10 = 2(x-10)
Solving for x
22 = x, steph's age now. Jess is 34. (22+12)
CHECKING our Answer*** 10 years ago...
24 = 2*12
1. Bart is 16 years older than Lisa.
Let x and (x+16)represent lisa's and Bart's ages NOW respectively
Question states*** In 3 years...
(x+16) + 3 = 3(x+3)
Solving for x
10 = 2x,
x = 5, Lisa's age now. Jess is 21. (5+16)
CHECKING our Answer*** In 3 years...
24 = 3*8